OpenStudy (anonymous):

Which expression describes the rule of this sequence? 7, 14, 21, 28, 35… 7 + n 7n 7n n+ 7 + 7

OpenStudy (anonymous):

7+n 7+7=14 14+7=21 and so on

OpenStudy (anonymous):

Look at the sequence below. What is the next term in the sequence? Fifty-one, forty-seven, forty-three, thirty-nine, thirty-five, and so on. thirty-four thirty-two thirty-one twenty-nine

OpenStudy (anonymous):

do 47-51=n (number) 35-n= answer

OpenStudy (anonymous):

Actually the first problem is \[ a_n= 7 n\\ a_1=7(1)=7\\ a_2=7(2)=14\\ a_3=7(3) =21\\ a_4=7(4) =28\\ \]

OpenStudy (anonymous):

I don't understand @RockStar02 it say what is the next sequence. not term..

OpenStudy (anonymous):

RIght. you have to figure the equation out. do 47-51=n (number) 35-n= answer

OpenStudy (anonymous):

31.

OpenStudy (anonymous):

Can you explain the rest to me ?

OpenStudy (anonymous):

@RockStar02